### v1

# Text-Speech Language Models with Improved Cross-Modal Transfer by Aligning Abstraction Levels

Abstract
Text-Speech Language Models (TSLMs)—language models trained to jointly process and generate text and speech—aim to enable cross-modal knowledge transfer to overcome the scaling limitations of unimodal speech LMs. The predominant approach to TSLM training expands the vocabulary of a pre-trained text LM by appending new embeddings and linear projections for speech, followed by fine-tuning on speech data. We hypothesize that this method limits cross-modal transfer by neglecting feature compositionality, preventing text-learned functions from being fully leveraged at appropriate abstraction levels. To address this, we propose augmenting vocabulary expansion with modules that better align abstraction levels across layers. Our models, SmolTolk, rival or surpass state-of-the-art TSLMs trained with orders of magnitude more compute. Representation analyses and improved multimodal performance suggest our method enhances cross-modal transfer.
Multimodal Language Models, Text-Speech Language Models, Speech Language Models, Language Models, Modality Alignment
1 Introduction
Generative auto-regressive language modeling has proven to be a powerful tool well beyond text. Since the paradigm-shifting arrival of Large Language Models (LLMs), a highly prolific line of research has focused on applying the same type of model to discrete representations (tokens) of other modalities, such as images (Chen et al., 2020), video (Yu et al., 2024), and audio (Lakhotia et al., 2021; Borsos et al., 2023). This unified approach has enabled the development of generative multimodal LMs, capable of jointly modeling different modalities (Aghajanyan et al., 2022; Chameleon Team, 2024; Rubenstein et al., 2023), paving the way for AIs that learn patterns across data of diverse nature and transfer knowledge between them (Aghajanyan et al., 2023). Given the remarkable capacities of text LLMs, a key goal is to enable cross-modal knowledge transfer between them and other modalities.
Because of its high mutual information with text, speech is well-suited for cross-modal transfer from text LMs. At the same time, there is a strong incentive to pursue this approach: unimodal speech LMs scale poorly, requiring orders of magnitude more compute to reach comparable linguistic competence (Cuervo & Marxer, 2024). This inefficiency suggests that simply increasing model and dataset sizes is unlikely to be viable, making cross-modal transfer not just promising but perhaps necessary. Accordingly, several works have explored joint text-speech modeling through Text-Speech Language Models (TSLMs) (Nguyen et al., 2024; Zhang et al., 2023; Défossez et al., 2024; Zeng et al., 2024), most following a common approach: expanding a pre-trained text LM’s vocabulary to accommodate speech by appending new embeddings and linear projections at the input and output respectively, then training on mixed speech and text data. While these methods demonstrate positive transfer, they remain heavily dependent on large amounts of paired text-speech data and still show a considerable performance gap between text and speech.
In this paper, we hypothesize that a key limitation of previous methods is their failure to account for the hierarchy of feature compositionality in text and speech. We note that: (1) Speech tokens, which encode acoustic and phonetic information, are at a lower level of abstraction than the word-like [^1] Assuming the widely used sub-word tokenization. tokens expected at the input and output of text LMs, as spoken words consist of multiple speech tokens. (2) LMs extract increasingly abstract features across layers, with the last few layers refining higher-level features into lower-level ones predictive of next tokens (Valeriani et al., 2023; Cheng et al., 2025). Thus, we argue that:
  1. 1.
    Since cross-modal transfer aims to enable functions learned during text pre-training to process speech’s linguistic information, speech tokens should first be composed into the word-like features these functions expect as input.
  2. 2.
    The output representations of text LMs are optimized to predict next words. Next-word information is useful for predicting next-speech tokens, and cross-modal transfer should leverage it. However, since next-speech tokens are at a lower level of abstraction than next words, the text LM’s output representations should be refined to predict next-speech tokens.
  3. 3.
    In speech language modeling, next-word information is most useful near word boundaries (or plausible boundaries), which are sparse relative to in-word tokens. Most of the time (within words), the information in the last layers for next-speech token prediction pertains to the current word and speech token—features that should be abstracted in the earlier layers. Therefore, we hypothesize that regular vocabulary expansion, which predicts speech tokens solely from text LM output representations, encourages the model to retain low-level information in later layers, possibly preventing it from abstracting higher-level features.
We posit that the linear functions used in vocabulary expansion methods are ill-suited for addressing points 1 and 2 and lack mechanisms to mitigate point 3. In this work, we test our hypotheses and aim to address these potential limitations. Our main contributions are:
  • We augment vocabulary expansion to better align abstraction levels across the model’s layers. To address points 1 and 2, we introduce additional speech-specific layers at the model’s input and output. To address point 3, we propose a learnable dynamic layer pooling mechanism that allows the speech output layers to adaptively switch between attending to low-level and next-word predictive representations.
  • We apply our method to the SmolLM family of language models (Ben Allal et al., 2024), which range in size from 135 million to 1.7 billion parameters. The resulting models, named SmolTolk, consistently outperform baselines trained with regular vocabulary expansion. Our largest model, SmolTolk-2B, rivals or surpasses orders-of-magnitude larger TSLMs.
  • We analyze the contributions of different architectural components through ablations and representation analyses. Our results suggest that each proposed element improves overall performance and contributes to facilitate cross-modal transfer.
2 Text-Speech Language Models
Text-Speech Language Models (TSLMs) model the joint probability of text and speech token sequences as
$P(\mathbf{w}=w_{1},\dots,w_{n})=\prod_{i=1}^{n}P(w_{i}|w_{1},\ldots,w_{i-1}),$ (1)
where $w_{i}\in\mathcal{V}_{t}\cup\mathcal{V}_{s}$, with $\mathcal{V}_{t}$ and $\mathcal{V}_{s}$ denoting text and speech vocabularies. TSLMs are typically decoder-only transformers (Vaswani et al., 2017) optimized to minimize the negative log-likelihood:
$\mathcal{L}_{LM}=-\sum_{i=1}^{n}P(w_{i}|w_{1},\ldots,w_{i-1}).$ (2)
Tokens are mapped to embeddings via a linear function $E\in\mathbb{R}^{(|\mathcal{V}_{t}|+|\mathcal{V}_{s}|)\times d}$, where $d$ is the embedding dimension. The sequence $E(w_{1}),\dots,E(w_{n})$ is processed by a stack of decoder-only transformer layers, producing contextual representations $(\mathbf{c}_{1},\dots,\mathbf{c}_{n})$, where $\mathbf{c}_{i}\in\mathbb{R}^{d}$ and each $\mathbf{c}_{i}$ depends on $\mathbf{c}_{\leq i}$. A linear projection $U\in\mathbb{R}^{d\times(|\mathcal{V}_{t}|+|\mathcal{V}_{s}|)}$ maps these to logits defining $P(w_{i+1}|\mathbf{c}_{i})$.
Text tokens are typically obtained via sub-word tokenization (Sennrich et al., 2016), while speech tokens are derived through quantization of self-supervised representations. Linguistic tokens [^2] Often referred to in the literature as ”semantic tokens”, though we argue this is a misnomer as they primarily encode phonetic information (Choi et al., 2024). are obtained from masked language modeling models like HuBERT (Hsu et al., 2021) and capture phonetic content. Acoustic tokens, extracted from autoencoder models, preserve speech signal details, including paralinguistic and acoustic variability (Zeghidour et al., 2021).
Training. TSLMs are trained via vocabulary expansion and speech fine-tuning of text LMs. Vocabulary expansion extends the embedding function and output projections over $\mathcal{V}_{t}$ to include $\mathcal{V}_{s}$, while the rest of the LM remains unchanged. Fine-tuning methods vary in data mixture: Rubenstein et al. (2023) train on mixed speech-text tasks (TTS, ASR, speech-to-speech translation), while Chou et al. (2023) use word-level alignments to switch modalities within a sequence. This interleaved text-speech strategy was shown to be crucial for cross-modal transfer, later validated and scaled up by Nguyen et al. (2024) and Zeng et al. (2024), achieving state-of-the-art speech LM performance.
Figure 1: General diagram of our proposed architecture. Input tokens are processed through modality-specific embedding tables. Speech representations (blue) pass through speech-specific layers, bridging the gap between speech and text inputs, before merging with text embeddings (green) in the text LM backbone. A weighted average of the backbone’s representations, computed using fixed and dynamic learned weights yields a representations better suited for speech prediction, which is processed through output speech-specific layers to predict the next speech tokens. In parallel, text tokens are predicted from the final text LM representation.
3 Our Method
Our method is illustrated in Figure 1. Below, we describe our model’s architecture and its training process.
3.1 Model
We apply the embedding function $E$ to the input sequence composed of text and speech tokens $\mathbf{w}$, yielding a sequence of text and speech embeddings $(E(w_{1}),\dots,E(w_{n}))=(\mathbf{z}_{1},\dots,\mathbf{z}_{n}$).
Input speech adapter. Contiguous chunks of speech embeddings are processed through an input adapter $A_{in}:\mathbb{R}^{d}\to\mathbb{R}^{d}$, a stack of decoder transformer layers. The input adapter is meant to compose the speech token embeddings into higher-level representations, which we believe should facilitate cross-modal transfer by matching the abstraction level expected by the text LM input. For a contiguous chunk of speech embeddings $(\mathbf{z}_{i},\dots,\mathbf{z}_{i+k})$, the input adapter outputs a sequence $(\mathbf{z}^{\prime}_{i},\dots,\mathbf{z}^{\prime}_{i+k})$.
The output of the input adapter and the text embeddings are fed into the text LM transformer at their respective positions in the input. For instance, given the input sequence $(w_{1},w_{2},w_{3},w_{4},w_{5})$, where only the third element is a text token, the sequence passed to the transformer layers after the adapter would be $(\mathbf{z}_{1},\mathbf{z}_{2},E(w_{3}),\mathbf{z}_{4},\mathbf{z}_{5})$. For each transformer layer $l$, we obtain a sequence of contextual representations $(\mathbf{c}_{1}^{(l)},\dots,\mathbf{c}_{n}^{(l)})$. As in regular text LMs and TSLMs, the text output logits are computed by applying $U$ to the contextual representations $\mathbf{c}_{i}^{(L)}$ at the last layer $L$. The speech output logits are computed as described next.
Dynamic layer pooling and speech input residual. We believe speech language modeling requires the model to switch between two modes of operation depending on whether a word is being generated or a new word is about to start. Within a word, the model should use low-level representations encoding the spoken word and the current speech token within that word, as these features fully determine the next speech token to be emitted. When generating a new word, the model should instead rely on representations predictive of upcoming words, such as those in the later layers of the text LM. To enable this behavior, we use a learnable mechanism that attends to different layers’ representations in an input-dependent manner. A linear layer selector $S:\mathbb{R}^{d}\times\mathbb{R}^{L}$ maps a contextual representation $\mathbf{c}_{i}^{\prime}$ to a vector of weights $\boldsymbol{\omega}_{i}=(\omega_{i}^{(1)},\dots,\omega_{i}^{(L)})$. These weights are softmax-normalized and used to compute an input-dependent weighted average yielding a contextual multi-level representation $\bar{\mathbf{c}}_{i}$:
$\boldsymbol{\omega}_{i}$ $=\text{Softmax}(S(\mathbf{c}_{i}^{\prime}))$ (3)
$\bar{\mathbf{c}}_{i}$ $=\sum_{l=1}^{L}\omega_{i}^{(l)}\mathbf{c}_{i}^{(l)}$
A key question is which representation to use as the layer selector’s input $\mathbf{c}_{i}^{\prime}$. We found that last layer’s representations $\mathbf{c}_{i}^{(L)}$ were not well suited as they often resulted in the selector collapsing to select a single layer. Rather than searching for the best layer—which would likely vary across different architectures—we use a weighted average of the contextual representations with learned input-independent weights:
$\mathbf{c}_{i}^{\prime}=\sum_{l=1}^{L}\phi^{(l)}\mathbf{c}_{i}^{(l)}$ (4)
where $\boldsymbol{\phi}=(\phi^{(1)},\dots,\phi^{(L)})$, $\phi^{(l)}\in\mathbb{R}$ are learned weights.
To provide information about the current speech token, we add a residual connection from the speech input embeddings to the multi-level contextual representation:
$\bar{\mathbf{c}}_{i}^{\prime}=\bar{\mathbf{c}}_{i}+\mathbf{z}_{i}$ (5)
$\bar{\mathbf{c}}_{i}^{\prime}$ contains both the information selected by layer pooling from the text LM layers and the current speech token.
Output speech adapter. The output adapter $A_{out}:\mathbb{R}^{d}\to\mathbb{R}^{d}$, a stack of decoder transformer layers, takes $\bar{\mathbf{c}}_{i}^{\prime}$ as input and refines it into a representation predictive of upcoming speech tokens, upon which the speech output logits are computed by applying $U$.
3.2 Training
We train our model to optimize Equation 2 on sequences from a data mixture similar to that proposed by Chou et al. (2023), including unimodal speech and text samples, as well as interleaved text-speech samples. Unlike previous works, we do not include ASR or TTS samples.
Two stage training. Training follows a two-stage process. In the first stage, the text LM backbone is frozen, and only the newly added modules are trained on interleaved text-speech modeling for approximately 3% of the total training iterations. This stage is introduced to mitigate text capability forgetting, as suggested by preliminary experiments. In the second stage, the full model is trained on the complete data mixture for the remaining iterations.
Preventing layer selector collapse. In larger models, we observed that the layer selector $S$ sometimes collapses early in training, attending to a single layer. To mitigate this, we add an entropy maximization term to the loss to encourage diversity in its output:
$\mathcal{L}=\mathcal{L}_{LM}+\beta\frac{1}{n}\sum_{i=1}^{n}\sum_{l=1}^{L}\omega_{i}^{(l)}\,\ln{(\omega_{i}^{(l)})}$ (6)
where $\beta$ is a hyperparameter that controls the entropy term.
4 Experimental setup
4.1 Models and training
We use SmolLM models (Ben Allal et al., 2024) as text LM backbones, available in three sizes: 135 million, 360 million, and 1.7 billion parameters. We refer to the models resulting from applying our method to the SmolLM backbones as SmolTolk-150M, SmolTolk-400M and SmolTolk-2B. Following Nguyen et al. (2024), and to improve the modeling of long-range dependencies, we increase the RoPE base frequency from 10,000 to 100,000.
The input and output adapters consist of transformer layers matching the text backbone architecture. In initial experiments, we observed that using more than two adapter layers led to only marginal or no improvements; thus, we use two layers for all our models. This aligns with findings from Turetzky & Adi (2024) on adapting speech representations to frozen text LMs. Across models, adding the adapters introduces a parameter overhead between 11% and 15%.
For speech tokenization, we follow Hassid et al. (2023), using the same tokenizer, which quantizes HuBERT representations extracted at 25 Hz into a 500-token vocabulary. As is common in speech LM training on linguistic tokens, we collapse adjacent token repetitions.
All LMs are optimized with AdamW (Loshchilov & Hutter, 2019) and a weight decay of 0.1. Models use a constant learning rate of 3e-4 for text and backbone parameters in the 135M, 360M, and 1.7B baselines, and 1e-4 for SmolTolk-2B. SmolTolk models apply a trapezoidal schedule for speech-specific parameters in stage 1 (1B tokens, 1000 steps), with a 100-step linear warmup and a maximum learning rate 10× that of text parameters, decaying over the final 20% of stage 1. For SmolTolk-400M and SmolTolk-2B, we set $\beta=0.01$ in Equation 6.
We use a batch size of 1 million tokens with the full 2048-token context. Each batch contains equal proportions of speech, text, and interleaved text-speech. All models are trained for 16 billion tokens, except the largest models, which are trained for up to 32 billion.
For more details on models and training see Appendix A.
4.2 Evaluation
[…]
We also report text performance on MMLU (Hendrycks et al., 2021) before (pre) and after (post) speech training to assess whether fine-tuning causes forgetting of text capabilities, as seen in other TSLMs (Nguyen et al., 2024; Défossez et al., 2024). We evaluate MMLU following the guidelines for the SmolLM models: https://huggingface.co/HuggingFaceFW/ablation-model-fineweb-edu#evaluation.
Baselines. We compare SmolTolk to models trained with the same text LM backbones and data but using regular vocabulary expansion, referring to these as baselines. We also compare against state-of-the-art TSLMs: Spirit LM (Nguyen et al., 2024), Moshi (Défossez et al., 2024), and the 1.5-billion and 9-billion models from Zeng et al. (2024).
Model Params. Tokens BLIMP tStoryCloze sStoryCloze MMLU
[…]
Previous TSLMs
Spirit LM (Nguyen et al., 2024) 7B $\sim$175B 73.3 59.7 95.8 90.5 78.6 94.3 74.0 66.3 64.7 71.7 37.7 / 39.0
Moshi (Défossez et al., 2024) 7.7B 2.1T 58.8 83.0 60.8 49.8 / 54.3
Zeng et al. (2024) 1.5B 1.5B 1T 77.5 81.4 90.1 55.4 58.6 64.0
Zeng et al. (2024) 9B 9B 1T 83.0 85.0 93.6 62.4 63.2 76.3
Ours
Baseline 1.7B 1.7B 16B 79.9 56.3 92.8 77.5 72.6 67.3 72.5 53.0 57.0 57.6 40.0 / 40.0
Baseline 1.7B 1.7B 32B 79.8 58.1 92.9 81.3 76.3 74.0 73.5 55.1 59.0 59.2 39.2 / 40.0
SmolTolk-150M 150M 16B 79.4 58.0 88.4 82.0 75.2 81.0 64.1 55.0 58.8 58.4 30.0 / 30.2
SmolTolk-400M 400M 16B 79.8 59.4 91.3 84.6 80.9 85.0 68.4 57.5 62.3 62.1 34.0 / 34.2
SmolTolk-2B 2B 16B 80.2 61.4 92.6 87.5 83.9 86.0 73.2 60.0 64.0 63.4 40.0 / 40.0
SmolTolk-2B 2B 32B 80.2 61.9 92.6 87.6 84.3 87.1 73.6 61.4 64.2 64.2 40.1 / 40.0
Table 1: Downstream evaluations. The best model in each task is shown in bold and underlined. The second best is shown in bold. For Spirit LM we report the results for the open-weights version. For other models we present the results reported by the authors.
4.3 Data
Speech datasets. We use a collection of publicly available English speech datasets for training: LibriSpeech (Panayotov et al., 2015), LibriLight (Kahn et al., 2020), SWC (Baumann et al., 2019), Tedlium (Hernandez et al., 2018), People’s Speech (Galvez et al., 2021), Vox Populi (Wang et al., 2021), and sTinyStories (Cuervo & Marxer, 2024). These datasets contain a total of 10.89 billion speech tokens.
Text datasets. We use a 12-billion-token subset of the SmolLM corpus (Ben Allal et al., 2024). Unlike Nguyen et al. (2024), we include math and code data, aiming to better preserve text capabilities. Our data distribution matches that used for pre-training SmolLM models, as reported in https://github.com/huggingface/smollm/blob/main/pre-training/.
Text-Speech datasets. We use the forced aligner from Pratap et al. (2024) to obtain word alignments for the LibriHeavy (Kang et al., 2024), sTinyStories, and SWC datasets. Interleaved samples are generated on the fly during batch sampling by randomly switching modalities within the input sequence. Following Nguyen et al. (2024), we randomly select word spans so that each text sequence contains 10–30 words and each speech sequence 5–15 words, balancing the proportion of speech and text tokens in each sample.
For more details on the datasets see Appendix B.
5 Experiments and results
Figure 2: Scaling of the LibriSpeech dev set negative log-likelihood (NLL) and tStoryCloze accuracy across modalities with respect to training compute (in FLOPs).
Table 1 presents the benchmark results. For brevity, we report only the results for the 1.7-billion baseline. Smaller baseline models underperformed relative to the larger one. An expanded results table is provided in Appendix D.
Our method significantly outperforms the baseline using regular vocabulary expansion across all tasks. One might attribute this difference to the parameter increase induced by the added modules. To better account for this factor, Figure 2 shows the scaling behavior of the negative log-likelihood (NLL) on the LibriSpeech dev set and the tStoryCloze benchmark as a function of compute, which helps adjust for model size differences and enables a fairer comparison. The figure demonstrates that, across the entire compute range, our models consistently outperform their respective baselines. Notably, SmolTolk-150M surpasses the 360-million baseline and rivals or surpasses the 1.7-billion baseline, suggesting that factors beyond model size drive the performance difference. Interestingly, in the baselines, cross-modal alignment scales only in the 1.7-billion parameter model. In contrast, our architecture demonstrates this capability even at smaller model sizes, indicating a more efficient approach to cross-modal learning. Scaling behavior for other metrics is presented in Appendix D.
Model tStoryCloze sStoryCloze
[…]
SmolTolk-400M 91.3 84.6 80.9 85.0 68.4 57.5 62.3 62.1
   –Dyn. pooling 90.8 84.0 80.1 83.9 68.2 57.5 60.9 61.6
   –Layer pooling 91.4 82.6 77.8 82.0 68.8 57.3 60.1 60.1
   –In Adapter 90.5 82.3 70.1 75.0 68.2 55.4 56.2 57.1
   –Out Adapter 90.7 80.7 76.1 84.1 68.1 54.9 60.0 60.0
   –Adapters 89.9 77.6 58.8 63.7 68.0 52.6 51.5 54.8
   –Residual 91.0 83.1 80.7 82.3 68.8 56.6 61.4 60.8
Baseline 360M 90.4 74.1 53.1 53.8 68.4 54.0 52.1 53.1
Table 2: Ablation Study. “–” denotes removal. ”–Dyn. pooling” uses fixed learned weights instead of dynamic ones from the layer selector, while ”–Layer pooling” entirely disables multi-layer pooling, relying only on the last text LM layer.
[…]
Our setup, including the baselines, exhibits less deterioration in text MMLU performance relative to other TSLMs. We attribute this to our decision to use a text fine-tuning distribution that matches the one used during pre-training.
5.1 Ablation study
Overall, the results indicate that our design choices enhance multimodal performance. To better understand each component’s contribution, we conduct an ablation study in Table 2 on the medium-size model by systematically removing elements and evaluating their impact.
We observe that removing any component degrades performance across most metrics, confirming the importance of our design choices. Eliminating all adapters results in the steepest drop—especially in cross-modal transfer, highlighting their role in bridging representations. The input adapter seems to be of greater importance for cross-modal transfer than the output adapter, underscoring the importance of merging modalities early in processing. Layer pooling also provide consistent gains, demonstrating the benefits of allowing the model to use multiple abstraction levels for speech language modeling. Removing the residual connection also causes a consistent decrease in performance.
Figure 3: Intrinsic dimensionality of representations across layers.
Figure 4: Fraction of variance explained by projecting one modality onto the other’s top 50 principal components across layers.
5.2 Representation analysis
To gain deeper insight into our design choices, we analyze learned representations across model variants. We investigate how the introduced architectural components influence feature abstraction and alignment between text and speech. We focus on two aspects: (1) the model’s capacity to abstract high-level features and (2) the shared structure of text and speech representations.
To assess (1), we follow Valeriani et al. (2023) and use the intrinsic dimensionality of representations as a proxy for compositionality. To evaluate (2), we compute the principal components of speech and text representations on paired data and measure how much variance is explained when projecting one modality’s representations onto the other’s principal components. The intuition is that knowledge transfer can be quantified by how much the model utilizes the same subspaces to represent equivalent data across modalities. We apply these analyses to the medium size models, including the architectural ablations from Table 2. For details on the experimental setup, see Appendix C. The results are shown in Figures 4 and 4.
From Figure 4, we observe that different components of our architecture significantly impact the model’s ability to compose higher-abstraction features. All our models, except the one without layer pooling, achieve a higher intrinsic dimension, suggesting that layer pooling is essential for enabling compositionality. The effect of the adapters is also evident: the absence of input adapters leads to lower compositionality in earlier layers, while the absence of output adapters results in an overall reduction in intrinsic dimension.
Figure 4 supports similar conclusions, showing that our architecture enables the highest degree of subspace overlap between modalities. As before, adapters and layer pooling are crucial for cross-modal transfer. The absence of input adapters leads to low subspace overlap in earlier layers, while the absence of output adapters reduces overlap in later layers. We hypothesize that this occurs because the model repurposes later layers to produce representations predictive of upcoming speech tokens. In this study the speech input residual has minimal effect.
What is dynamic layer pooling learning?. Figure 5 shows the weights assigned by the layer selector $S$ across layers (top) and for the last layer—specialized in next token prediction—(bottom) across a speech input, alongside word boundaries. The layer weights follow a pattern consistent with our hypothesis, switching between low-level representations and those predictive of next words. Notably, spikes in last-layer weights often align with word boundaries, suggesting the model leverages next-word predictive representations at these points. To validate this, we compute a speech word segmentation score using a peak detector on the last layer’s attention weights as a word boundary predictor (see Appendix C for details). Table 3 compares our approach with SCPC (Bhati et al., 2022), a state-of-the-art unsupervised speech segmentation model, on the TIMIT test split (Garofolo et al., 1993). Our method outperforms SCPC, supporting our hypothesis on dynamic pooling behavior. However, we do not claim our segmentation method is unsupervised, as the model is trained on interleaved data that implicitly provides word boundary information. Interestingly, segmentation performance scales inversely with model size. We hypothesize that this may be because the greater capacity of output adapters in larger models allows them to rely less on backbone features. For segmentation examples of other sequences and models see Appendix D.
Figure 5: Selector $S$ layer weights across a speech input sequence for SmolTolk-2B. Vertical bars indicate word boundaries.
Model Precision Recall F-1 R-val
SCPC (Bhati et al., 2022) 30.3 20.3 24.5 40.5
SmolTolk-150M 50.5 46.7 48.4 56.8
SmolTolk-400M 37.6 37.2 37.4 46.8
SmolTolk-2B 33.3 29.4 31.2 43.4
Table 3: Word segmentation scores on TIMIT test set using dynamic pooling last layer weights as word boundary predictors.
6 Related work
Chou et al. (2023), Spirit LM (Nguyen et al., 2024), and Zeng et al. (2024) train TSLMs by jointly modeling text, speech, and interleaved text-speech without using text as an intermediate representation for speech generation. We adopt a similar training approach but propose a different speech modeling architecture that enhances multimodal performance. Moreover, compared to Spirit LM and Zeng et al. (2024), we use smaller, less diverse open speech datasets and far fewer computational resources while achieving comparable or, at times, superior performance.
Other state-of-the-art TSLMs follow a hierarchical process, first generating text as an intermediate representation to condition speech generation, akin to TTS. Notable works include SpeechGPT (Zhang et al., 2023), Moshi (Défossez et al., 2024), and Llama-Omni (Fang et al., 2024). Unlike these, our model does not rely on text generation before speech synthesis. Llama-Omni is particularly relevant, as it employs input speech adapters with temporal downsampling and an MLP layer, and an output adapter with temporal upsampling and transformer decoder layers. In contrast, our adapters avoid temporal downsampling/upsampling, eliminating latency, and use transformer decoder layers in the input adapter as well, as we have evidence from Valeriani et al. (2023); Cheng et al. (2025) that these layers can perform the composition processes we posit are necessary.
A closely related work architecturally is LAST (Turetzky & Adi, 2024), where adapters with the same architecture as ours feed and generate speech from a text LM, primarily to learn a speech tokenization from HuBERT representations. Unlike LAST, we apply multi-layer pooling before the output adapter, perform multimodal training, target cross-modal capabilities, and achieve state-of-the-art performance.
7 Discussion and future work
7.1 Implications
On speech language modeling. We believe our results make a strong case for future TSLM research to adopt our proposed architecture. Our method can complement other approaches to improving speech language modeling, such as large-scale diverse datasets (Zeng et al., 2024), additional losses (Lin et al., 2024), or coarser speech tokenization (Baade et al., 2025). The latter may better interface with the text LM and reduce the load—and thus the required capacity—of adapters.
Our dynamic multi-layer pooling method may also be relevant for textless speech LM research, as a similar representation hierarchy likely emerges in speech LMs, reflecting speech’s inherent feature hierarchy. Thus, speech predictive layers should adaptively switch between attending to low- and high-level representations to facilitate high-level feature abstraction, as suggested by the experiments in Section 5.2.
On interfacing other modalities to text LMs. Overall, our results support our initial hypotheses, highlighting the importance of accounting for feature compositionality when adapting text LMs for speech modeling. We believe this insight may extend to other modalities, as multimodal generative LMs use similar vocabulary expansion strategies, e.g., in image modeling (Wang et al., 2022; Aghajanyan et al., 2022; Chameleon Team, 2024). Although models like Chameleon (Chameleon Team, 2024) tout the absence of modality-specific modules as a strength, our findings suggest this may come at the cost of cross-modal alignment.
More broadly, we think of text as the most compressed representation humans use to interact with the world, closest to a Platonic representation of abstract concepts, following the thought-provoking idea from Huh et al. (2024). As such, we believe most other perceptual modalities naturally exist at lower abstraction levels (i.e., less compressed) and require composition at the input and refinement at the output to properly interface with text LMs. Therefore, we suggest that applying the ideas used in this work for speech to other modalities is a worthy direction for future research.
7.2 Limitations
Direct applications. Since our model is not instruction fine-tuned, its applicability to common speech LM tasks, such as spoken conversational agents, is limited. Future work could explore instruction tuning and broader evaluations. Additionally, using linguistic tokens makes our model less expressive—one of speech LMs’ key advantages over cascaded approaches combining text LMs with ASR and TTS. Future work could apply our architecture to model more expressive speech representations.
Scale. Our models are small by current standards, and the observed phenomena may differ at larger scale. Figure 2 shows that cross-modal transfer scaling tends to improve in larger models without our architecture. This suggests that simpler models could match our performance at scale, echoing The Bitter Lesson (Sutton, 2019). However, in line with The Bitter Lesson, our design imposes no additional priors, allowing the model to learn to effectively revert to the simpler baseline architecture if optimal. Therefore, we do not expect it to bottleneck performance. Instead, given the observed gains, it is likely to remain beneficial.
Novelty. While perhaps not explicitly stated, the hypotheses studied in this work clearly underlie many prior multimodal LM studies using adapters, as they are fairly intuitive and perhaps even obvious. However, previous works doing vocabulary expansion gave little consideration to these aspects, and we show this may have significantly harmed performance. Moreover, our ablation and representation analyses strengthen the understanding of the underlying phenomena. Additionally, to the best of our knowledge, we are the first to argue for and demonstrate the usefulness of multi-layer representations in speech language modeling.
8 Conclusions
We introduced a more effective approach to text-speech language modeling by enhancing vocabulary expansion with speech-specific adapters and dynamic layer pooling. These components improve abstraction alignment across model layers, enabling more effective cross-modal transfer. Our evaluations and representation analyses show that this method better integrates speech into text LMs.
Our SmolTolk models achieve state-of-the-art performance using far fewer computational resources and smaller, less diverse datasets than previous works. Notably, SmolTolk-2B rivals or surpasses much larger TSLMs. These results underscore the importance of hierarchical feature compositionality in multimodal learning. Beyond speech, our findings suggest that aligning feature abstraction levels may be key to adapting text LMs for other modalities.
Author contributions
Santiago proposed and led the project, defining hypotheses, model and experimental design, and developed most of the training, model and analysis codebase. Adel played a key role in data preprocessing and synthesis, text and cross-modal evaluations, and model inference, while also contributing insights on model design. He maintained the codebase with Santiago’s assistance. Ricard provided supervision throughout and contributed significantly to the data pipeline, experimental codebase, and analysis. Yanis assisted with experiment implementation, execution, and discussions. Other authors contributed to early-stage discussions and supervision during the JSALT 2024 workshop.
Acknowledgements
Santiago and Ricard are grateful to the French National Research Agency for their support through the ANR-20-CE23-0012-01 (MIM) grant, as well as to the Institute of Convergence ILCB, which is supported by grants from France 2030 (ANR-16-CONV-0002) and the Excellence Initiative of Aix-Marseille University (A*MIDEX). Adel is funded by the Cambridge Trust and is grateful to his former supervisor in Avignon, professor Yannick Estève, for involving him in this project and providing sponsorship. Other team members have also received funding from the European Union’s Horizon 2020 research and innovation program under the Marie Skłodowska-Curie grant agreement No. 101007666.
This work was supported by HPC GENCI-IDRIS under grants AD011014044R1, A0161014876, AD011015344, A0161015099, AD011013061R3, and A0161014871.
Finally, we extend our gratitude to the JSALT 2024 workshop and the Center for Language and Speech Processing at Johns Hopkins University for generously hosting the event, fostering the emergence of fruitful collaborations and friendships in a vibrant and stimulating environment.
Appendix A Models details
Table 4 describes the SmolLM (Ben Allal et al., 2024) backbones architectural hyperparameters. All SmolTolk models use two layer input and output adapters with the same architecture as the backbone layers. Table 5 describes the resulting models after vocabulary expansion.
Model Num. Layers Num. Heads Num. KV Heads Emb. Dim. Hidden Dim.
SmolLM-135M 30 9 3 576 1536
SmolLM-360M 32 15 5 960 2560
SmolLM-1.7B 24 32 32 2048 8192
Table 4: Backbones architectural hyperparameters.
Model Backbone Num. Layers Text params Speech Params
Baseline-135M SmolLM-135M 30 135M 0.29M
Baseline-360M SmolLM-360M 32 360M 0.48M
Baseline-1.7B SmolLM-1.7B 24 1.7B 1M
SmolTolk-150M SmolLM-135M 34 135M 15M
SmolTolk-400M SmolLM-360M 36 360M 40M
SmolTolk-2B SmolLM-1.7B 28 1.7B 270M
Table 5: Models description.
As mentioned in Section 3.1, we use a linear layer with bias as the dynamic layer selector $S$. We experimented with simple non-linear MLP selectors; however, these were prone to collapse and resulted in worse overall performance. That said, a more carefully designed non-linear selector could potentially perform better. We also explored alternative ways to define the contextual representation $\mathbf{c}_{i}^{\prime}$ (Equation 3) used as input for the layer selector. Instead of a learned weighted average, we tried concatenating low-dimensional linear projections from each layer’s representations, but this performed worse.
Regarding training, we tuned the learning rate for each model, including baselines, to be as high as possible without causing instabilities or increasing text data validation loss, which we considered a sign of text capability forgetting. We also experimented with learning rate schedules for text/backbone parameters, but a constant rate performed better.
For our experiments, we used NVIDIA H100 GPU nodes, each featuring four NVIDIA H100 80GB SXM5 GPUs, dual Intel Sapphire Rapids 48-core processors, 512GB of RAM, and four NVIDIA ConnectX-7 400Gb/s InfiniBand network adapters. SmolTolk-2B used a per-GPU batch size of 8 million tokens and was trained on 64 GPUs across 16 nodes, taking approximately 10 hours to process 32 billion tokens. All models were trained using bfloat16 mixed precision with FlashAttention-2 (Dao, 2024) and PyTorch compile.
Appendix B Datasets details
Table 6 presents the size of the datasets used for each modality and their respective sampling ratios. In each batch, each modality was sampled with equal probability. We experimented with reducing the proportion of text samples per batch to as low as 10%, as their primary purpose is merely to preserve text capabilities. However, we observed that this significantly impacted text downstream performance. The sampling ratios of text datasets were set according to the pre-training setup of SmolLM models, as specified in https://github.com/huggingface/smollm/blob/main/pre-training/. Speech datasets were sampled based on their size—that is, uniformly across tokens.
For interleaved text-speech datasets, we upsampled the Spoken Wikipedia Corpora (SWC), as it is a knowledge-dense dataset, and we hypothesized that this would enhance downstream performance on knowledge-based tasks. Additionally, we expected it to better align with the distribution of article-style samples in the text pre-training data, potentially facilitating cross-modal transfer. However, we did not explicitly evaluate these capabilities in this work. The sampling ratios for LibriHeavy and sTinyStories were defined in proportion to their size.
Modality Dataset Tokens Sampling ratio
Text Speech
Text FineWeb-Edu (Ben Allal et al., 2024) 4B 0.7
Cosmopedia-v2 (Ben Allal et al., 2024) 4B 0.15
Python-Edu (Ben Allal et al., 2024) 2B 0.08
FineMath (Liu et al., 2024) 2B 0.06
Speech LibriSpeech (Panayotov et al., 2015) 67M (960 hours)
LibriLight (Kahn et al., 2020) 3.7B (53k hours)
SWC (Baumann et al., 2019) 32M (1k hours)
Tedlium (Hernandez et al., 2018) 0.1B (1.6k hours)
People (Galvez et al., 2021) 0.5B (7k hours)
Vox Populi (Wang et al., 2021) 1.6B (24k hours)
sTinyStories (Cuervo & Marxer, 2024) 4.8B (72k hours)
Interleaved text-speech LibriHeavy (Kang et al., 2024) 313M 3.1B (50k hours) 0.37
sTinyStories (Cuervo & Marxer, 2024) 800M 4.8B (72k hours) 0.53
SWC (Baumann et al., 2019) 3.6M 26M (800 hours) 0.1
Table 6: Datasets statistics. Speech datasets were sampled according to their size.
Appendix C Representation analyses setup
C.1 Intrinsic dimensionality and subspace overlap
The intrinsic dimensionality and subspace overlap are estimated using five batches of 10k representations each, totaling 50k samples. We compute each metric per batch and report average and deviations. To obtain each representation, we extract random subsequences of 20 words and use the final representation in the sequence as the sequence representation. For the intrinsic dimensionality text samples are randomly drawn from FineWeb-Edu, while speech samples are taken from the full set of speech datasets. For the subspace overlap we use paired samples from the sTinyStories dataset.
Since transformer architectures exhibit large activation outliers, we truncate feature elements (i.e., individual activations) that exceed the 95th percentile across the entire 50k sample set.
To estimate the intrinsic dimensionality we use the Generalized Ratios Intrinsic Dimension Estimator (GRIDE) (Denti et al., 2022) implementation in dadapy (Glielmo et al., 2022) and follow the procedure described by Cheng et al. (2025).
C.2 Word segmentation
We apply a peak detector to the sequences of last-layer dynamic weights, $\boldsymbol{\omega}^{(L)}$, using SciPy’s find_peaks tool. Performance is evaluated as a binary prediction task, where a peak indicates the prediction of a boundary at a given position. We allow a tolerance of one token to account for noise in the boundary annotations. As in standard binary prediction tasks, we use recall, precision, and the F1-score as performance metrics. Additionally, we use the R-value (Räsänen et al., 2009), which penalizes trivial over-segmentation solutions. To optimize performance, we tune the prominence parameter of the peak detector over a grid $(0,0.15]$ with steps of $0.01$ so as to maximize the R-value, following Kreuk et al. (2020).
Appendix D Additional results
Extended benchmark results, including the smaller baselines and other speech LMs from the literature, are presented in Table 7. Figure 6 illustrates the scaling of several metrics for the baseline and SmolTolk models.
Figures 7, 8, and 9 provide additional results on the analysis of multi-layer pooling for SmolTolk-2B, SmolTolk-400M, and SmolTolk-150M, respectively.
Model Params. Tokens BLIMP tStoryCloze sStoryCloze MMLU
[…]
Textless Speech LMs
GSLM (Lakhotia et al., 2021) 100M Ø 54.2 Ø 66.6 Ø Ø Ø 53.3 Ø Ø Ø
AudioLM (Borsos et al., 2023) 150M Ø 64.7 Ø Ø Ø Ø Ø Ø Ø
Hassid et al. (2023) cold-init 1.3B 1.3B 10.8B Ø 56.5 Ø Ø Ø Ø Ø Ø Ø
TWIST (Hassid et al., 2023) 1.3B 1.3B 10.8B Ø 57.0 Ø 70.6 Ø Ø Ø 52.4 Ø Ø Ø
TWIST (Hassid et al., 2023) 7B 7B 36B Ø 59.0 Ø 74.1 Ø Ø Ø 55.3 Ø Ø Ø
TWIST (Hassid et al., 2023) 13B 13B 36B Ø 59.2 Ø 76.4 Ø Ø Ø 55.4 Ø Ø Ø
Cuervo & Marxer (2024) best 823M 82B Ø 61.3 Ø 78.0 Ø Ø Ø 56.7 Ø Ø Ø
SyllableLM (Baade et al., 2025) 300M 1.2B Ø 63.7 Ø 75.4 Ø Ø Ø Ø Ø Ø
AlignSLM (Lin et al., 2024) 7B 7B Ø 62.3 Ø 86.8 Ø Ø Ø 61.1 Ø Ø Ø
Slam (scaled) (Maimon et al., 2025) 358M 16.7B Ø 61.1 Ø 84.2 Ø Ø Ø 61.3 Ø Ø Ø
Previous Text-Speech LMs
Spirit LM (Nguyen et al., 2024) 7B $\sim$175B 73.3 59.7 95.8 90.5 78.6 94.3 74.0 66.3 64.7 71.7 37.7 / 39.0
LAST (Turetzky & Adi, 2024) $\sim$390M 56.8
Moshi (Défossez et al., 2024) 7.7B 2.1T 58.8 83.0 60.8 49.8 / 54.3
Zeng et al. (2024) 1.5B 1.5B 1T 77.5 81.4 90.1 55.4 58.6 64.0
Zeng et al. (2024) 9B 9B 1T 83.0 85.0 93.6 62.4 63.2 76.3
Ours
Baseline 135M 135M 16B 79.0 52.0 87.0 73.2 53.3 52.7 63.9 54.0 53.8 53.7 30.3 / 30.2
Baseline 360M 360M 16B 79.8 52.4 90.4 74.1 53.1 53.8 68.4 54.0 52.1 53.1 34.5 / 34.0
Baseline 1.7B 1.7B 16B 79.9 56.3 92.8 77.5 72.6 67.3 72.5 53.0 57.0 57.6 40.0 / 40.0
Baseline 1.7B 1.7B 32B 79.8 58.1 92.9 81.3 76.3 74.0 73.5 55.1 59.0 59.2 39.2 / 40.0
SmolTolk-150M 150M 16B 79.4 58.0 88.4 82.0 75.2 81.0 64.1 55.0 58.8 58.4 30.0 / 30.2
SmolTolk-400M 400M 16B 79.8 59.4 91.3 84.6 80.9 85.0 68.4 57.5 62.3 62.1 34.0 / 34.2
SmolTolk-2B 2B 16B 80.2 61.4 92.6 87.5 83.9 86.0 73.2 60.0 64.0 63.4 40.0 / 40.0
SmolTolk-2B 2B 32B 80.2 61.9 92.6 87.6 84.3 87.1 73.6 61.4 64.2 64.2 40.1 / 40.0
Table 7: Downstream evaluations. The best model in each task is shown in bold and underlined. The second best is shown in bold. For Spirit LM we report the results for the open-weights version. For other models we present the results reported by the authors.
Figure 6: Scaling of the LibriSpeech dev set negative log-likelihood (NLL), R-val word segmentation score, and tStoryCloze and sStoryCloze accuracy across modalities with respect to training compute (in FLOPs).
Figure 7: Selector $S$ dynamic layer weights $\boldsymbol{\omega}$ across several speech input sequences for SmolTolk-2B. Vertical bars indicate word boundaries (Top and middle). Non-input dependent learned weights $\boldsymbol{\phi}$ (bottom-left). Distribution of dynamic weights $\boldsymbol{\omega}$ across layers (bottom-right).
Figure 8: Selector $S$ dynamic layer weights $\boldsymbol{\omega}$ across several speech input sequences for SmolTolk-400M. Vertical bars indicate word boundaries (Top and middle). Non-input dependent learned weights $\boldsymbol{\phi}$ (bottom-left). Distribution of dynamic weights $\boldsymbol{\omega}$ across layers (bottom-right).
Figure 9: Selector $S$ dynamic layer weights $\boldsymbol{\omega}$ across several speech input sequences for SmolTolk-150M. Vertical bars indicate word boundaries (Top and middle). Non-input dependent learned weights $\boldsymbol{\phi}$ (bottom-left). Distribution of dynamic weights $\boldsymbol{\omega}$ across layers (bottom-right).